Sprinkles285:

whats the square root 56?

Sprinkles285:

Helpppp

jhonyy9:

use this 8*7 = 56 8 = 2^3 = 2^2 *2 sqrt56 = sqrt(8*7) = (sqrt8 ) * (sqrt7) = (sqrt(2^2 *2) * (sqrt7) = (2sqrt2 )*(sqrt7) = ?